Part B: Please review the material below and answer the questions A woman, age 35, had suffered from anorexia nervosa for 10 years and was admitted to the hospital for treatment Her weight was about 60% ofthe ideal for her height (5 fet, 5 inches, 726 lbs). She was fed by gavage (ao gastric feeding) with 3L/day of a nutrient solution that provided 1000Cal/L and had the following composition Protein hydrolysate Fat Carbohydrate Sodium Potassium Chloride Magnesium Phosphate Calcium Zinc Manganese 0.37 moles N 42 g 125 g 21.7 mmole 2 mmole 28 mmole 8.3 mmole 16.1 mmole 15 mmole 157 mmole 45 mmole 15.7 mmole 0.6 mmole 161 mmole Copper lodine Iron Complete vitamin mixture No complications resulted from the treatment&on the basis of the positive N balance &the positive balances of several other nutrient minerals it was concluded that the mutriests were being used for the syathesis of nomal body mass. After 21 days, the patient increased 22 lbs in body weight 11. What is the protein equirvalent of the nutrient uid? What could be concluded concerning its biclogical value 12. What differences might be noted in the nature of the autrients for garage feeding as compazed to total parenteral feeding (hyperalimentation) 13. In the light of present knowledge what additional component) might be added to the diet to ensure otimu nutrition
Answer #1

11. What is the protein equivalent of the nutrient fluid? What could be concluded concerning its biological value.

Protein equivalent of the nutrient fluid:-
As given, Protein hydrolysate-0.37 moles N protein in eating regimen; It implies that 0.37*14=5.18 gof nitrogen. (moles=9/MW or 9= mol* MW) the esteem is looking great to Increase positive nitrogen balance in body.
Presently we can state the organic estimation of given protein is great.
Note:- Positive nitrogen balance implies increment the nitrogen maintenance in the body which will build the amino acids level and protein. This procedure is for working out. The development harmones and other body developing atoms are likewise working positive nitrogen balance.

12. What differences might be noted in the nature of the nutrients for gavage feeding as compared to total parenteral feeding (hyperalimentation)
Nature of nutrients for gavage feeding compared to total parental feeding (hyperalimentation):-
The gavage bolstering is greatly improved than the parental liquid (No uncertainty parental noneed to ingestion and processing) since we can utilize parental feed as most extreme farthest point however gavage nourishing is vastly improved. It is boundless normal marvels , which is more valuable to enhance body weight.

13. In the light of present knowledge what additional component(s) might be added to the diet to ensure optimum nutrition.
Additional components to be added to the diet to ensure optimum nutrition :-
The arrangement can build the creation of amino acids/Protein. She has a noteworthy issue of weight reduction. Egg whites contain diet can be included. Magnesium and zinc can likewise be added to enhance the nitrogen balance.

Please do liked and rate me...........